# Quantum coherence and the invisible Universe: Subradiance as a dark matter mechanism

### v3

@@ Appendix B The master equation
The master equation used for the analysis presented in Sec. 2.2 is as follows
@@ C.1 Initial Bloch angle and magnetization
$\left(\hat{R}^{-}\right)^{p}$ |$\Ket{r,m}=\left[\left(r+m\right)\ldots\left(r+m-p\right)\right.$
@@
$\Bra{r,m}\left(\hat{R}^{+}\right)^{p}$ $\left(\hat{R}^{-}\right)^{q}\Ket{r,m}=\left(r+m\right)\ldots\left(r+m-p\right)$
@@
At a given position $z$ the magnetization will be defined by the atoms in the vicinity (see equation 58) and we can approximately replace $\hat{R}^{\pm}$ by $\hat{R}^{\pm}_{k}$ in equation (61).
@@
$\Braket{r,m}{\left(\hat{R}^{+}_{k}\right)^{p}\left(\hat{R}^{-}_{k}\right)^{q}}{r,m}\simeq\delta_{pq}\,p!\left(2r\right)^{p},$ (62)
@@
$\Braket{r,m}{\left(\hat{R}^{+}_{k}\right)^{p}\left(\hat{R}^{-}_{k}\right)^{q}}{r,m}\simeq\delta_{pq}\left(r^{2}-m^{2}\right)^{p}.$ (63)
@@
$\Braket{\overline{r},\overline{m}}{\left(\hat{R}^{+}_{k}\right)^{p}\left(\hat{R}^{-}_{k}\right)^{q}}{\overline{r},\overline{m}}\simeq\delta_{pq}\left(\frac{N}{2}\right)^{p}.$ (66)
@@
$\Braket{\overline{r},\overline{m}}{\hat{R}^{+}_{k}\left(0\right)}{\overline{r},\overline{m}}\equiv\beta e^{i\varphi}$ (67)
@@ C.2 Linear regime solutions
$\left.+\frac{\alpha}{2}e^{-\tau/T_{2}}\left[e^{-\alpha z/2}\star J_{0}\left(2\sqrt{\frac{z\tau}{LT_{R}}}\right)\right]\right\}$ (76)
@@
The evolution time-scale of the system subjected to an incident field can be calculated as follows. As a first step, expanding the Bessel function for $z\tau\ll LT_{R}$ in equation (76) allows us to approximate
$B_{1}^{+}\left(z,\tau\right)\simeq B_{0}$ $\left[e^{-\alpha z/2}\left(1-e^{-\tau/T_{2}}e^{-2\sqrt{z\tau/LT_{R}}}\right)\right.$
$\left.+e^{-\tau/T_{2}}e^{-2\sqrt{z\tau/LT_{R}}}\right].$ (78)
If we further restrict ourselves to $T_{2}\ll T_{R}$, then expanding the square root about $zT_{2}/LT_{R}$ in equation (78) yields
$e^{-\tau/T_{2}}e^{-2\sqrt{z\tau/LT_{R}}}\approx e^{-\tau\left(1+\sqrt{zT_{2}/LT_{R}}\right)/T_{2}}$ (79)
and an evolution time-scale given by
$T_{\mathrm{tr}}\approx T_{2}\left(1+\sqrt{\frac{zT_{2}}{LT_{R}}}\right)^{-1}.$ (80)
@@ C.3 Non-coherent radiation intensity
where $\phi_{\mathrm{D}}=\lambda^{2}/A$ is the diffraction solid-angle at the end of our (cylindrical) system.
@@ Appendix D Collision cross-section
[…]
If two atoms enter into an elastic collision, say the first atoms of both pairs, then under the Born approximation each term will bring a scattering amplitude of the type [38]
$f\left(\mathbf{k}_{\mathrm{in}},\mathbf{k}_{\mathrm{out}}\right)=-\frac{m_{\mathrm{r}}}{2\pi\hbar^{2}}\int d^{3}r^{\prime}e^{-i\mathbf{k}_{\mathrm{out}}\cdot\mathbf{r}^{\prime}}V\left(\mathbf{r}^{\prime}\right)e^{i\mathbf{k}_{\mathrm{in}}\cdot\mathbf{r}^{\prime}}$ (88)
with $\mathbf{k}_{\mathrm{in}}=\left(\mathbf{P}_{B}-\mathbf{P}_{A}\right)/\hbar$, $\mathbf{k}_{\mathrm{out}}$ the scattered wave vector and $m_{\mathrm{r}}$ the reduced mass. For high relative speed between the colliding atoms, i.e., when $\left|\mathbf{P}_{B}-\mathbf{P}_{A}\right|^{2}/2m_{\mathrm{r}}\gg\hbar\omega_{0}$, it is expected that the scattering amplitude will be independent of their internal states. It therefore follows from equation (87) that under these circumstances the collision cross-section $\sigma\left(\mathbf{k}_{\mathrm{in}},\mathbf{k}_{\mathrm{out}}\right)=0$ since it consists of the square of the sum of the four corresponding scattering amplitudes.
Although such calculations need to be generalized, we find that at high relative speed two darkened entangled systems are likely to pass through each other, unimpeded, during a collision. They would not feel the presence of each other even though individual atoms may actually collide. This is due to the destructive interference resulting from the entanglement within each of the systems.

### v4

@@ Appendix B The master equation
The master equation used for the analysis presented in Sec. 2.2 is as follows [20, 22]
@@ C.1 Initial Bloch angle and magnetization
$\left(\hat{R}^{-}\right)^{p}$ |$\Ket{r,m}=\left[\left(r+m\right)\ldots\left(r+m+1-p\right)\right.$
@@
$\Bra{r,m}\left(\hat{R}^{+}\right)^{p}$ $\left(\hat{R}^{-}\right)^{q}\Ket{r,m}=\left(r+m\right)\ldots\left(r+m+1-p\right)$
@@
When considering states |$\Ket{r,m}_{\theta}$, such as those used in Sec. 2.1 for the two-atom case, it can be shown that a similar relation holds if we replace $\hat{R}^{\pm}$ by $\hat{R}^{\pm}_{k}$ in equation (61). That is, we write
${}_{\theta}\Bra{r,m}\left(\hat{R}^{+}_{k}\right)^{p}$ $\left(\hat{R}^{-}_{k}\right)^{q}\Ket{r,m}_{\theta}=\left(r+m\right)\ldots\left(r+m+1-p\right)$
@@
${}_{\theta}\Bra{r,m}\left(\hat{R}^{+}_{k}\right)^{p}\left(\hat{R}^{-}_{k}\right)^{q}\Ket{r,m}_{\theta}\simeq\delta_{pq}\,p!\left(2r\right)^{p},$ (63)
@@
${}_{\theta}\Bra{r,m}\left(\hat{R}^{+}_{k}\right)^{p}\left(\hat{R}^{-}_{k}\right)^{q}\Ket{r,m}_{\theta}\simeq\delta_{pq}\left(r^{2}-m^{2}\right)^{p}.$ (64)
@@
${}_{\theta}\Bra{\overline{r},\overline{m}}\left(\hat{R}^{+}_{k}\right)^{p}\left(\hat{R}^{-}_{k}\right)^{q}\Ket{\overline{r},\overline{m}}_{\theta}\simeq\delta_{pq}\left(\frac{N}{2}\right)^{p}.$ (67)
@@
${}_{\theta}\Bra{\overline{r},\overline{m}}\hat{R}^{+}_{k}\left(0\right)\Ket{\overline{r},\overline{m}}_{\theta}\equiv\beta e^{i\varphi}$ (68)
@@ C.2 Linear regime solutions
To determine the evolution time-scale of the system, $T_{\mathrm{tr}}$, when subjected to an incident field, we expand the Bessel function for $z\tau\ll LT_{R}$ in equation (77) to approximate
[…]
and we find
$T_{\mathrm{tr}}\approx T_{2}\left(1+\frac{zT_{2}}{LT_{R}}\right)^{-1}.$ (80)
That is, $T_{\mathrm{tr}}\simeq T_{2}$ for $T_{2}\ll T_{R}$ and $T_{\mathrm{tr}}\simeq T_{R}$ for $T_{2}\gg T_{R}$ at $z=L$.
When responding to initial internal fluctuations, a similar process yields
$B_{2}^{+}\left(z,\tau\right)$ $\approx-i\frac{\hbar\theta_{0}}{2\mu T_{R}}\frac{z}{L}e^{-\tau/T_{2}}e^{-z\tau/2LT_{R}}$ (81)
when $z\tau\ll LT_{R}$ in equation (78). We thus have
$T_{\mathrm{tr}}\approx T_{2}\left(1+\frac{zT_{2}}{2LT_{R}}\right)^{-1}$ (82)
with $T_{\mathrm{tr}}\simeq T_{2}$ for $T_{2}\ll T_{R}$ and $T_{\mathrm{tr}}\simeq 2T_{R}$ for $T_{2}\gg T_{R}$ at $z=L$.
@@ C.3 Non-coherent radiation intensity
where $\phi_{\mathrm{D}}=\lambda^{2}/A$ is the diffraction solid-angle at the end of our (cylindrical) system. We also note that equation (83) can be transformed such that
$I_{\mathrm{nc}}=B_{\nu_{0}}\left(T_{\mathrm{kin}}\right)\frac{\phi_{\mathrm{D}}}{2T_{R}},$ (84)
where $B_{\nu_{0}}\left(T_{\mathrm{kin}}\right)$ is the Planck function at frequency $\nu_{0}=\omega_{0}/2\pi$ and temperature $T_{\mathrm{kin}}$. The non-coherent intensity is therefore that of a blackbody into a solid angle $\phi_{\mathrm{D}}$ and spectral bandwidth $\left(2T_{R}\right)^{-1}$ (see equation 82).
@@ Appendix D Collision cross-section
For high relative speed between colliding atoms, i.e., when $\left|\mathbf{P}_{B}-\mathbf{P}_{A}\right|^{2}/2m_{\mathrm{r}}\gg\hbar\omega_{0}$ ($m_{\mathrm{r}}$ is the reduced mass), it is expected that the scattering amplitude will be independent of their internal states (see section D.1 below). If two atoms enter into a collision, say the first atoms of both pairs, then under the Born approximation each term will bring a scattering amplitude of the type [38, 39]
$f\left(\mathbf{k}_{\mathrm{in}},\mathbf{k}_{\mathrm{out}}\right)=-\frac{m_{\mathrm{r}}}{2\pi\hbar^{2}}\int d^{3}r^{\prime}e^{-i\mathbf{k}_{\mathrm{out}}\cdot\mathbf{r}^{\prime}}V_{1}\left(\mathbf{r}^{\prime}\right)e^{i\mathbf{k}_{\mathrm{in}}\cdot\mathbf{r}^{\prime}}$ (91)
with $\mathbf{k}_{\mathrm{in}}=\left(\mathbf{P}_{B}-\mathbf{P}_{A}\right)/\hbar$, $\mathbf{k}_{\mathrm{out}}$ the scattered wave vector and $V_{1}\left(\mathbf{r}\right)$ the component of the atomic interaction potential independent of the internal states. It therefore follows from equation (90) that under these circumstances the collision cross-section $\sigma\left(\mathbf{k}_{\mathrm{in}},\mathbf{k}_{\mathrm{out}}\right)=0$ since it consists of the square of the sum of the four corresponding scattering amplitudes [^3] An exchange term should be added to equation (90) to account for the symmetrization process stemming from the indistinguishability of the two identical atoms during the collision [38, 39]. For simplicity, we only consider the direct term since the same result (i.e., the vanishing of the corresponding scattering amplitude) is obtained for the exchange term.. We note that the vanishing of the scattering amplitude rests entirely on the (anti-symmetric) form of the entangled states, as in equation (89). This outcome would therefore hold for any other types of high-energy scattering processes (e.g., involving cosmic rays, etc.).
Although such calculations need to be generalized, we find that at high relative speed two darkened entangled systems are likely to pass through each other, unimpeded, during a collision. They would not feel the presence of each other even though individual atoms may collide. This is due to the destructive interference resulting from the entanglement within each of the systems.
@@
D.1 Independence on internal states for high energy scattering
In the center-of-mass of the system, the Hamiltonian for the scattering of two hydrogen atoms in the electronic ground state can be expressed as [40]
$\hat{H}=\frac{\hat{p}^{2}}{2m_{\mathrm{r}}}+\hat{V}_{1}\left(\mathbf{r}\right)+\hat{H}_{\mathrm{hf}}$ (92)
with $\mathbf{r}$ and $\mathbf{p}$ the relative position and (half of) the relative linear momentum between the two atoms, respectively. The hyperfine Hamiltonian results from the intra-atomic interactions due to the Fermi contact term [38], which dominates over any other interatomic spin coupling term, and is thus given by
$\hat{H}_{\mathrm{hf}}=\mathcal{A}\left(\hat{\mathbf{I}}_{1}\cdot\hat{\mathbf{S}}_{1}+\hat{\mathbf{I}}_{2}\cdot\hat{\mathbf{S}}_{2}\right),$ (93)
where $\hat{\mathbf{I}}_{j}$ and $\hat{\mathbf{S}}_{j}$ are the proton and electron spins for atom $j$, and $\mathcal{A}=\omega_{0}/\hbar$ a constant. In equation (92) (and 91), the interatomic potential $\hat{V}_{1}\left(\mathbf{r}\right)$ is independent of the internal states (and spins).
From a classical standpoint, the turning point $r_{\mathrm{min}}$ for a collision at energy $E_{\mathrm{kin}}=p^{2}/2m_{\mathrm{r}}$ will be approximately determined by $V_{1}\left(r_{\mathrm{min}}\right)\sim E_{\mathrm{kin}}$ [41]. In cases such as the bullet cluster [32, 27], with collision speeds $\sim 10^{3}\,\mathrm{km\,s^{-1}}$, we find that $E_{\mathrm{kin}}\sim 10^{9}\hbar\omega_{0}$ and therefore $V_{1}\left(r_{\mathrm{min}}\right)\gg\hbar\omega_{0}$, or equivalently $V_{1}\left(r_{\mathrm{min}}\right)\gg H_{\mathrm{hf}}$. It follows that the hyperfine Hamiltonian is completely negligible in the region where the scattering occurs.
Since the internal states are determined by the atomic spins, we observe a decoupling between the internal and external degrees of freedom on account of the large separation in the corresponding energy scales. In other words, the scattering amplitude is largely independent of the atoms’ internal state and can be evaluated with equation (91) (including the hyperfine interaction brings a correction term of the order of $\hbar\omega_{0}/E_{\mathrm{kin}}$).